On January 1, 2017, Wood Corporation leases a piece of equipment from Prior Corporation and properly accounts for the equipment as a finance lease. Under the agreement, Wood will make 7 annual payments of $550,000 each January 1st. At the end of 7 years, Wood has the option of buying the equipment for $200,000, when the estimated fair value will be $400,000. If Wood's incremental borrowing rate is 7%, what is the present value of the minimum lease payments?

$3,088,659.12

$2,754,865.26

$3,296,146.76

None of the above
Answer #1

$3,296,146.76

Present value of minimum lease payments = PV of annual lease payments + PV of residual value

PV of annual lease payments

Using financial calculator , SET BGN mode
Input: PMT = 550000

N = 7

I/Y = 7

Solve for PVas

3,171,596.81

PV of residual value

Using financial calculator
Input:FV = 200000

N = 7

I/Y = 7

Solve for PV as

124,549.95

Total PV =3171,596.81+ 124,549.95= $3,296,146.76